### v1

@@ 4.3 Multimodal Datasets
ChatGPT 51.57 - - -2.81 -6.28 32.80 41.63 74.85 37.07 76.27 77.51 48.93 28.86 31.44 34.20 192.25 - -
LLaMA1-7B -50.97 - - -47.87 -49.57 -35.96 -22.66 -20.17 -25.81 -52.30 -37.98 -41.46 -46.55 -11.85 -22.24 -93.47 - -
LLaMA1-13B -51.84 - - -17.08 -11.25 -35.95 -19.93 -16.56 -12.02 -41.17 -22.56 -15.42 -19.51 -19.05 -32.17 -59.44 - -
LLaMA2-7B -37.93 - - -11.24 -8.86 -24.30 -31.27 -39.26 -9.84 -55.93 -38.14 -31.74 -36.70 -27.36 -18.75 -88.05 - -
LLaMA2-13B -47.79 - - -22.48 -21.48 -14.94 -20.75 -38.36 -26.87 -60.73 -35.68 -34.86 -44.32 -29.70 -28.18 -71.98 - -
Mixtral-AWQ 27.18 - - -3.26 -9.31 16.91 11.78 42.48 3.21 58.90 12.68 14.72 -4.69 20.96 28.54 142.48 - -
Gemma 13.00 - - 36.04 24.82 24.39 18.49 25.77 4.31 -12.56 10.11 19.31 9.50 21.96 10.43 120.24 - -
Flan-T5-XXL 63.30 - - 27.27 27.54 46.64 42.11 66.43 45.81 114.81 49.09 50.64 35.73 33.23 53.63 193.09 - -
Gemini-V 65.79 41.38 39.89 43.90 43.01 32.63 33.11 67.56 42.40 49.36 35.59 42.45 41.77 45.96 3.97 264.42 2 7
OpenFlamingo -58.69 -51.74 -47.50 -53.19 -48.84 -51.10 -35.87 -58.67 -40.51 -66.59 -52.68 -53.01 -51.20 -41.85 -47.47 -90.70 0 0
Fromage -69.73 -74.55 -85.98 -74.51 -85.75 -81.65 -70.07 -82.98 -74.15 -89.57 -83.14 -80.18 -75.18 -59.26 -54.39 -99.37 0 0
LLaVA-v0-7B -9.43 21.19 10.22 24.45 13.70 -43.60 -0.55 -0.51 3.98 -27.15 -24.06 -11.13 9.98 -55.75 -29.25 -79.68 0 0
LLaVA-v0-13B -3.41 33.74 28.05 37.15 31.96 -5.42 -0.48 -14.73 7.15 -8.21 -4.02 9.93 34.94 -33.90 -28.01 -76.11 0 0
LLaVA-v1.6-7B 36.14 13.17 27.05 14.98 30.02 22.74 28.39 6.49 34.04 -18.06 36.98 30.46 35.70 33.68 23.77 -15.09 0 0
LLaVA-v1.6-13B 48.30 20.65 16.24 23.36 19.47 46.72 35.57 38.27 22.98 72.55 46.66 37.21 36.85 40.21 31.31 55.97 0 1
MiniGPT4 13.93 43.05 23.04 46.26 26.33 45.01 17.41 24.29 26.56 16.10 7.62 19.08 31.61 24.48 8.88 -85.01 0 2
mPLUG-Owl -38.73 -17.61 -28.86 -16.13 -27.33 -30.05 -29.66 -30.29 -32.37 -49.21 -40.77 -39.54 -28.60 -3.56 -18.94 -74.00 0 0
mPLUG-Owl2.1 34.49 -5.04 14.97 -3.37 17.62 37.02 23.82 32.78 16.45 34.78 41.75 28.37 27.19 24.90 19.26 18.30 0 0
AdapterV2 20.91 42.27 41.71 45.14 45.48 2.61 -1.44 4.00 13.86 -24.76 -7.94 7.58 17.19 -9.36 -0.69 -71.76 0 4
VPGTrans -28.38 12.87 9.32 16.18 13.36 11.08 -15.41 -22.78 -9.53 -18.54 -27.94 -28.40 -10.38 -32.33 5.67 -90.78 0 0
MultiGPT -60.50 -52.95 -47.57 -52.77 -46.87 -51.39 -46.94 -49.45 -49.25 -73.79 -43.68 -46.95 -38.31 -28.71 -34.72 -91.09 0 0
LaVIN-7B -7.49 -43.67 -42.30 -42.17 -40.59 -54.40 -28.57 -38.72 -40.57 -51.82 -46.66 -50.47 -28.06 -6.67 7.94 -76.79 0 0
LaVIN-13B -2.81 -32.20 -45.98 -31.05 -44.38 -30.76 -8.13 -21.66 -13.81 -16.86 -40.28 -36.02 -16.56 -16.80 -23.97 -53.80 0 0
Lynx -51.03 -0.21 5.68 2.46 9.72 -21.40 -26.20 -62.03 -18.58 -72.15 -8.90 -4.54 4.67 -20.54 -39.09 -63.09 0 0
Fuyu-8B -21.47 -51.31 -33.77 -50.78 -32.50 18.30 14.21 -33.21 6.53 61.31 57.52 19.12 -19.56 22.49 15.60 -95.28 0 1
LaVIT -16.95 6.43 11.11 8.47 14.45 -7.85 -6.05 -5.16 -7.59 -43.01 -12.02 1.43 -2.22 6.60 -10.28 -98.07 0 0
Qwen-VL-Chat 32.27 17.06 26.06 20.02 29.53 41.36 13.89 22.51 3.97 78.03 61.58 45.57 22.21 20.21 30.63 45.97 0 1
BLIP-2 72.46 34.00 35.98 36.55 39.55 51.44 42.25 68.21 47.21 112.19 46.54 49.29 40.76 34.71 55.81 212.59 2 10
InstructBLIP 77.79 43.46 42.63 46.49 46.46 59.12 41.30 60.89 45.35 108.13 42.82 49.63 44.89 35.86 58.52 228.26 8 13
5 Experimental Results and Analysis
We evaluate the zero-shot performance of 31 LLMs (as detailed in Table 2) across 16 datasets (listed in Table 3), utilizing 10 distinct instructions per model (as shown in Figures 3, 4, and 5). The related results in a total of 4960 (31 $\times$ 16 $\times$ 10) experimental results, each yielding an accuracy score denoted as $acc_{mdi}$ from Eq. 1. Although most of the evaluated datasets could rely on text context for inference, PuzzleVQA and MMMU are exceptions, necessitating multimodal contexts. For ScienceQA, which primarily depend on the image modality to answer questions, we also explore the potential to extract knowledge from models using solely the image modality. Our comprehensive evaluation incorporates a range of metrics, including the Best Performance metric, the Mean Relative Gain metric, the Stability metric, and the Adaptability metric. Detailed analyses of these metrics and insightful conclusions related to them are presented in the subsequent sections.
5.1 Best Performance
We present the best performance ($A^{\tilde{i}}$), calculated using Eq. 2, achieved by different models across various multimodal datasets. The metric serves as benchmarks for the performance of different models on each dataset. Based on these results, we make the following observations:
5.1.1 Performance of Various Models on Full Datasets
We first present the best performance of various models across full datasets, as detailed in Table 4.
(1) Comparison between open-source models and closed-source models: While the Flan-T5 series, including Flan-T5, BLIP-2, and Instruct-BLIP, are strong open-source contenders, closed-source models such as Gemini-V typically outperform their open-source counterparts in multimodal reasoning tasks involving vision-text contexts. Notably, Gemini-V achieves the best performance across the most datasets; ChatGPT also delivers excellent performance using text-only contexts. However, it remains uncertain whether these closed-source models have been specifically trained on these datasets, a factor that could contribute to their superior performance. Another possible explanation for the observed discrepancies is that these closed-source models may be significantly larger than the open-source models evaluated. Unfortunately, we do not have specific information about the scale of the closed-source models. (2) Flan-T5 backbone vs. LLaMA backbone in the open-source models: Within open-source models, InstructBLIP stands out by securing top performance on three datasets and clinching the first position in the “Total” column, with BLIP-2 closely following in second place. Notably, both InstructBLIP and BLIP-2 utilize Flan-T5-XXL as their backbone model. Flan-T5-XXL is recognized as the best-performing large language model among the 8 LLMs evaluated (as shown in the first part of Table 4). It highlights the exceptional performance capabilities of Flan-T5-XXL. The superior performance of Flan-T5-XXL could be attributed to its Encoder-Decoder architecture, unlike the Decoder-only architecture of the LLaMA series models. The encoder module in Flan-T5-XXL offers significant advantages in multimodal representation learning, which is crucial for handling multimodal reasoning tasks with complex multimodal contexts. (3) LLMs vs. MLLMs: For the cumulative best performance of models across all datasets—excluding PuzzleVQA and MMMU, as indicated by the “Total” column—we observe that LLMs exhibit superior performance compared to certain MLLMs, such as Openflamingo, Fromage, LaVIT, and others. The phenomenon may be due to theses MLLMs tend to specialize in traditional visual language (VL) capabilities, including language generation, recognition, knowledge reasoning, spatial reasoning, and OCR, linked to the image modality. Meanwhile, these models might not sufficiently prioritize capabilities like multimodal alignment, multimodal fusion, and etc. Such capabilities are critical for effectively handling multimodal reasoning tasks that involve both vision and text contexts. (4) Different scale checkpoints: In the same model series, larger versions, evidenced by increased parameters such as in LLaMA-7B vs. LLaMA-13B, LLaVA-v0-7B vs. LLaVA-v0-13B, and LaVIN-7B vs. LaVIN-13B, generally outperform their smaller counterparts across most datasets. It suggests that models with more parameters and extensive training are likely to achieve superior performance. Furthermore, the latest model iterations, such as the mPLUG-OWL and LLaVA series, exhibit significant enhancements over previous versions, indicating that advancements in the novel architecture and more training substantially boost model efficacy. (5) Various instructions: On one hand, performance peaks vary within the same model when applied to different datasets, contingent on the instructions used. On the other hand, across various models, the optimal performance on a specific dataset often requires model-specific instructions. Despite its importance, the area of selecting the most suitable instructions for diverse datasets and models is under-researched and represents a promising avenue for future exploration. (6) More challenge tasks: Closed-source models like Gemini-V, significantly outperform open-source models on more challenging datasets, including MOSI-7 and MOSEI-7 which involve multiple classification categories, MNRE with long-tail effect, and more complex reasoning datasets such as PuzzleVQA and MMMU. There remains considerable potential for improvement in open-source models addressing these tasks. (7) With/Without text context on ScienceQA: ScienceQA primarily relies on vision context to answer questions. Accordingly, we conduct comparative experiments that exclude textual context, as detailed in the “Text context” part of Figure 3. The results of these experiments are shown in Figure 6. We observe that while most MLLMs demonstrate better performance on the ScienceQA dataset when textual context is included, the degree of improvement is limited. It suggests that the presence or absence of textual context has a limited impact on the performance of the ScienceQA dataset.
5.1.2 Full datasets vs. Subsets
Given the high cost and limited availability of GPT-4V and Claude3-opus, we draw inspiration from MultiPoint (Yang et al., 2023c) and GMP (Yang et al., 2023b) to tailor our approach. Specifically, we construct subsets by randomly sampling 10% of the data from the corresponding complete test datasets, ensuring that the distribution remains consistent with that of the original datasets. We then exclusively assess GPT-4V and Claude3-V on these subsets across all tasks, selecting corresponding experimental results from other models for comparison. The comparative analysis, focusing on the best performance metric between the full dataset and its subset, is depicted in Figure 7.
We employ the “LLMs-Total” metric to sum the best accuracy across all datasets, excluding PuzzleVQA and MMMU, to gauge overall performance for each LLM. Similarly, the “MLLMs-Total” metric sums the best performance across all datasets including PuzzleVQA and MMMU for assessing MLLM performance. The subplots “LLMs-Total” and “MLLMs-Total” indicate minimal performance differences between models on full datasets and their subsets, suggesting that our evaluation of models like GPT-4V and Claude3 on subsets is an accurate reflection of their overall performance across tasks. Moreover, GPT-4V consistently achieves superior performance on most tasks. It excels particularly in challenging datasets with extensive classification categories, such as MOSI-7 and MOSEI-7, datasets featuring long-tail effects like MNRE, and complex reasoning tasks including PuzzleVQA and MMMU, where it significantly outperforms other models. Notably, some open-source models, including InstructBLIP and BLIP2, also deliver competitive results on many subsets. An additional observation is that larger datasets tend to show less variation in experimental outcomes. The stability underscores the value of using subsets from large datasets, as they tend to provide more representative results.
@@ 5.1.1 Performance of Various Models on Full Datasets
BLIP-2 5.73 0.99 1.29 5.46 7.86 3.05 0.31 3.62 0.42 3.43 9.04 3.55 0.53 0.69 1.52 1.89 2
InstructBLIP 0.73 1.46 1.07 4.15 4.44 0.62 0.62 5.02 0.51 4.24 7.02 3.29 0.92 0.65 1.27 3.90 6

### v2

@@ 5.1.1 Performance of Various Models on Full Datasets
ChatGPT 61.41 - - 2.21 1.33 33.34 37.53 58.55 34.77 74.52 82.43 52.06 45.26 28.13 30.68 79.58 - -
LLaMA1-7B -48.40 - - -48.88 -48.63 -37.10 -25.66 -28.66 -28.45 -53.58 -44.28 -46.89 -46.30 -17.31 -26.08 -95.75 - -
LLaMA1-13B -51.18 - - -18.16 -12.29 -39.41 -24.07 -25.46 -14.96 -42.92 -25.19 -18.97 -16.05 -23.89 -35.26 -75.26 - -
LLaMA2-7B -35.38 - - -6.07 -2.89 -26.21 -33.90 -46.06 -11.52 -57.28 -43.07 -37.23 -36.11 -33.83 -22.16 -92.47 - -
LLaMA2-13B -47.46 - - -23.44 -23.19 -17.15 -25.69 -44.34 -29.69 -61.47 -40.84 -39.93 -44.08 -35.05 -32.39 -82.78 - -
LLaMA3.1-8B-Instruct -7.66 - - 8.39 7.63 -2.91 12.07 -6.98 2.63 -14.61 0.28 6.81 21.30 15.18 2.52 97.95 - -
Mixtral-AWQ 34.42 - - 3.88 0.19 16.84 6.79 27.87 1.14 56.47 7.64 8.89 0 18.73 25.29 50.85 - -
Gemma 20.33 - - 47.36 38.92 25.60 14.90 13.58 2.21 -13.75 12.68 22.13 23.09 17.54 7.43 35.41 - -
Flan-T5-XXL 73.93 - - 34.72 38.92 47.00 38.02 51.50 43.86 112.63 50.35 53.13 52.89 30.50 49.99 82.47 - -
Gemini-V 29.10 -20.14 14.17 40.74 10.52 29.98 24.20 45.47 50.42 -3.20 -3.42 21.42 34.58 66.87 22.32 94.82 1 3
OpenFlamingo -66.71 -53.38 -57.18 -51.64 -58.45 -51.11 -38.71 -63.69 -36.50 -78.28 -63.44 -57.15 -51.35 -33.70 -36.56 -95.11 0 0
Fromage -75.47 -72.43 -68.42 -74.13 -88.49 -81.48 -70.79 -84.53 -71.88 -92.83 -87.20 -82.27 -75.26 -52.95 -45.58 -99.66 0 0
LLaVA-v0-7B -29.85 -46.48 -41.62 20.76 -12.51 -45.06 -6.98 -13.52 10.12 -52.97 -45.30 -24.01 3.93 -49.10 -17.35 -89.40 0 0
LLaVA-v0-13B -25.15 -67.29 -65.52 33.36 1.49 -7.22 -6.94 -26.41 13.22 -40.87 -31.14 -5.83 27.92 -24.21 -15.23 -87.26 0 0
LLaVA-v1.6-7B 5.38 -17.94 -12.19 12.46 0.29 20.65 19.31 -8.63 41.41 -47.10 -3.13 10.89 28.34 52.32 45.68 -54.26 0 0
LLaVA-v1.6-13B 14.82 -20.13 2.57 19.99 -8.33 43.62 25.98 19.65 29.76 10.89 4.41 16.59 29.30 59.82 54.56 -15.91 0 3
MiniGPT4 -11.39 -29.48 -22.40 41.99 -2.92 42.23 8.69 6.28 33.21 -24.96 -23.26 1.55 24.31 41.81 28.09 -92.14 0 1
mPLUG-Owl -51.63 -39.29 -43.82 -17.88 -43.61 -31.29 -33.59 -39.34 -27.84 -66.64 -56.56 -47.32 -31.18 9.72 -4.55 -86.22 0 0
mPLUG-Owl2.1 4.02 -24.51 -13.88 -5.57 -9.34 34.21 15.04 14.65 22.82 -13.04 1.67 9.54 20.66 42.36 40.18 -36.67 0 0
AdapterV2 -6.59 -46.57 -55.69 41.21 11.45 0.09 -8.55 -9.48 20.20 -51.28 -34.77 -8.92 10.58 4.85 16.99 -85.20 0 1
VPGTrans -43.92 -62.55 -53.01 12.25 -13.51 8.77 -20.50 -33.39 -4.39 -47.00 -47.45 -38.11 -14.56 -23.58 24.73 -95.06 0 0
MultiGPT -68.31 -71.87 -65.42 -52.14 -57.55 -51.68 -48.83 -55.39 -45.53 -82.97 -56.97 -52.24 -38.90 -18.68 -21.09 -95.15 0 0
LaVIN-7B -26.60 -55.69 -57.57 -43.74 -54.21 -54.99 -32.75 -46.14 -36.92 -68.54 -61.57 -57.54 -32.05 7.00 27.58 -87.28 0 0
LaVIN-13B -23.10 -41.64 -51.60 -32.30 -57.07 -31.38 -14.52 -33.05 -9.08 -46.96 -57.00 -45.79 -21.40 -4.87 -9.61 -74.77 0 0
Lynx -61.34 -68.47 -71.54 -0.42 -16.14 -22.35 -31.13 -67.05 -14.45 -81.81 -34.28 -17.50 0.44 -9.58 -27.47 -80.63 0 0
Fuyu-8B -38.57 -28.36 -39.52 -51.43 -47.75 16.44 6.51 -42.37 12.75 4.92 12.95 1.90 -22.48 39.66 36.38 -97.33 0 0
LaVIT -35.15 -33.32 -44.78 5.81 -12.38 -9.29 -12.69 -17.83 -2.38 -63.04 -37.07 -13.85 -7.72 21.83 5.26 -98.99 0 0
LLaMA-3.2-11B-Vision 17.78 61.75 54.90 -5.58 27.47 -11.14 4.00 -0.50 -13.49 59.37 51.33 12.86 -17.12 -30.99 -44.78 132.87 1 1
MiniCPM-V2.6 65.91 108.09 95.10 11.02 50.05 24.88 20.67 48.38 -25.38 63.65 60.17 40.08 10.27 -3.69 -31.43 112.60 0 6
BLIP-2 33.45 -20.81 -4.40 33.24 7.23 48.37 32.10 44.64 55.24 36.48 5.64 27.44 33.17 53.54 83.36 67.07 2 5
InstructBLIP 37.34 -21.23 -8.80 42.63 12.41 55.90 31.24 38.49 53.27 33.86 2.30 27.47 37.10 54.82 86.55 75.62 4 7
GLM-4V-9B 72.26 105.34 95.99 4.41 43.21 -4.55 13.65 39.65 -21.05 62.74 64.72 36.75 -10.50 -22.38 -25.30 122.47 2 4
InternVL2.5-8B 62.12 95.95 80.16 2.47 44.40 17.59 14.42 46.45 -0.46 83.17 59.67 38.17 14.49 6.08 -32.77 111.1 0 3
DeepSeek-VL2-tiny 22.98 113.17 99.40 -8.99 23.03 -4.78 17.68 -28.24 -32.96 48.85 52.23 13.51 3.69 -93.03 -35.12 63.04 2 2
DeepSeek-VL2-small 20.95 48.05 40.61 -21.39 25.14 -32.48 -0.43 23.19 -21.50 42.03 44.40 14.35 -0.74 -93.87 -44.01 36.85 0 0
DeepSeek-VL2 5.64 93.44 67.98 -2.47 42.39 7.92 17.57 16.49 -22.74 57.75 40.14 19.65 -0.81 -32.36 -34.57 78.20 0 0
Qwen-VL-Chat 2.50 -27.45 -9.92 16.52 -0.21 38.59 5.99 5.20 9.87 15.57 14.69 23.83 15.74 37.05 53.65 -21.78 0 0
Qwen2-VL-7B 45.94 56.13 61.85 7.08 40.20 -3.61 14.69 56.41 4.45 77.74 59.60 44.92 8.53 -12.44 -25.40 132.87 2 3
Qwen2-VL-72B-Instruct 27.47 51.71 51.85 5.39 50.20 42.12 18.23 68.24 16.94 95.60 42.69 34.13 26.97 35.99 -26.68 120.77 3 3
Qwen2.5-VL-3B-Instruct 39.48 53.59 42.88 6.55 44.73 11.91 10.11 48.67 4.13 80.97 61.76 23.03 -9.13 -36.63 -21.74 118.28 0 2
Qwen2.5-VL-7B-Instruct 56.65 81.83 79.81 9.79 48.28 -0.85 26.35 47.72 8.72 87.90 64.21 32.46 3.18 8.32 -26.08 126.24 0 5
@@
LLaVA-v1.6-13B 4.70 0.51 1.51 5.17 5.48 2.61 1.85 2.59 5.79 3.07 5.95 2.86 4.42 1.43 2.76 5.74 0
MiniGPT4 6.71 2.56 2.35 4.99 13.06 2.48 9.17 6.62 8.02 7.10 3.93 4.23 3.91 1.65 5.95 0.69 1
AdapterV2 2.64 7.78 7.88 10.65 8.46 6.68 23.57 11.08 21.00 7.26 7.14 12.06 20.19 15.25 6.61 2.13 0
LLaMA-3.2-11B-Vision 15.55 6.72 7.87 6.87 11.39 5.39 7.86 4.34 6.02 4.25 2.36 10.04 6.11 4.75 4.43 2.39 0
MiniCPM-V2.6 2.65 1.28 1.28 0.85 0.65 3.45 1.96 3.17 2.02 1.02 0.78 1.07 3.42 3.03 0.87 1.44 2
BLIP-2 5.73 0.99 1.29 5.46 7.86 3.05 0.31 3.62 0.42 3.43 9.04 3.55 0.53 0.69 1.52 1.89 3
InstructBLIP 0.73 1.46 1.07 4.15 4.44 0.62 0.62 5.02 0.51 4.24 7.02 3.29 0.92 0.65 1.27 3.90 4
GLM-4V-9B 3.46 3.15 1.60 3.32 3.36 8.62 4.01 5.91 4.39 1.69 0.45 1.57 6.41 2.52 0.51 1.69 1
InterVL2.5-8B 4.63 1.98 1.63 1.27 1.08 1.65 2.70 1.90 2.72 1.08 0.51 0.96 0.95 3.07 0.91 3.07 0
DeepSeekVL2-tiny 5.54 2.00 1.40 2.80 8.52 1.86 3.07 3.35 6.06 2.30 2.37 1.34 6.89 5.14 0.74 1.95 0
Qwen2-VL-7B-Instruct 3.15 5.36 6.44 0.87 7.67 6.18 8.16 1.41 4.95 2.26 0.50 0.89 2.50 2.29 1.45 2.39 1
Qwen2-VL-72B-Instruct 2.31 2.74 2.49 1.39 1.14 1.22 4.25 1.18 2.68 0.49 3.67 1.01 3.75 1.43 1.34 2.40 2
Qwen2.5-VL-3B-Instruct 2.68 9.42 8.70 1.60 1.56 2.41 3.50 7.98 7.32 2.14 0.56 6.51 5.02 2.91 0.56 1.65 0
Qwen2.5-VL-7B-Instruct 2.21 4.52 2.00 1.82 2.06 5.27 0.86 7.51 6.37 3.26 0.31 4.80 4.86 2.53 1.44 1.80 1
5.2 Mean Relative Gain
5.2.1 Mean Relative Gain of Models
To assess the comprehensive performance of each model, we calculate the mean relative gain for each model across all instructions on each dataset, as detailed in Eq. 3. The results are tabulated in Table 5. The metrics “Wins1” and “Wins3” denote the number of tasks where a model or an instruction achieves the highest and top-three aggregated performances, respectively. Similar to the best model performance metric, models equipped with the Flan-T5-XXL backbone, such as InstructBLIP, BLIP-2, and Flan-T5-XXL itself, exhibit strong performance. The latest models, notably Qwen2.5-VL and MiciCPM-V2.6, demonstrate superior performance across all instructions, particularly in VQAMC tasks. However, their strong performance should be interpreted with caution, as they may have been pre-trained on the ScienceQA dataset, potentially introducing data leakage that gives them an advantage over other models. Closed-source models also show excellent results.
5.2.2 Mean Relative Gain of Instructions
We continue to evaluate the overall efficacy of each instruction by calculating the mean relative gain (MRGI) across all models, as detailed in Eq. 4. The outcomes are presented in Table 6. Due to the different evaluation datasets for LLMs and MLLMs—with LLMs excluding PuzzleVQA and MMMU and MLLMs including these datasets—we compute MRGI separately for each group. Our analysis reveals that Instructions # 1, # 2, # 5, and # 7 generally perform better, with their detailed designs illustrated in Figures 3 and 4. Notably, Instructions # 5 and # 7, which evolve from simple “Question-Answer” format of # 2, show superior performance. The effectiveness likely stems from the prior training of models on related QA tasks, where specific terms like “Question” and “Answer” enhance model responsiveness. Furthermore, the inclusion of option terms in Instructions # 5 and # 7 appears to influence performance on specific datasets positively. We find an interesting phenomenon, where models performed better on VQAMC tasks when given Instruction # 5, which includes options, while Instruction # 2, without options, yielded superior results for MSA tasks. This pattern may be attributed to the distinct nature of these tasks. In VQAMC, each question has unique answers, making the provided options valuable hints for response generation. Conversely, in MSA tasks, the label space remains constant across examples, rendering the inclusion of options less beneficial.
5.3 Stability
5.3.1 Stability of Models